OpenStudy (anonymous):

Solve the equation 2^4x=4^x-6 help me please ;<

OpenStudy (anonymous):

2^4x=2^2(x-6) 4x=2x-12 2x=-12 X=-6

OpenStudy (campbell_st):

well rewrite the right hand side as a power of 2 \[2^{4x} = (2^2)^{x - 6}\] and after using the power of a power rule gives \[2^{4x} = 2^{2x - 12}\] now you have the same base... equate the powers and solve for x \[4x = 2x - 12\]

OpenStudy (anonymous):

Thanks alot <33